OpenStudy (anonymous):

. Convert each of the following: a. 100 in. =________ft. b. 400 yd.=________ ft. c. 300 ft.=_________ in. d. 372 in.=________ yd.

OpenStudy (unklerhaukus):

12 in = 1 ft 1 in = 1/12 ft 100 in = 100/12 ft =

OpenStudy (unklerhaukus):

1 yd = 3 ft